# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 5.2 Transportation
Transportation serves as the arteries of a city, acting as bridges that connect different urban entities Doi (2015). As a result, the comprehensive utilization of multi-modal data focuses a significant proportion of attention on downstream transportation-related tasks Miller et al. (2016); Bwire and Zengo (2020); Sinha et al. (2019); Li et al. (2023a).
@@ 5.2.1 Public Transportation Service
The advancement of a city is primarily demonstrated by enhancements in its transportation systems Doi (2015); Guo et al. (2020); Dai et al. (2015). Public transportation, in comparison to private transportation, is recognized for its superior environmental sustainability and energy efficiency Miller et al. (2016). Guided by government supervision, public transportation embodies the features of public goods. In addition to fundamental passenger transportation, there are also freight transportation services, along with emerging services such as food delivery and ride-hailing, all of which contribute to enhancing urban convenience. Different from Zheng et al. (2014a), where transportation systems were categorized based on the type of vehicles used, this classification of services is based on their specific application scenarios. Public transportation applications can primarily be categorized into three main groups: Safety, Flow Control, and Efficiency.
@@
Safety is the most fundamental and urgent requirement. Consequently, roads, being high-risk areas for public transportation safety, have prompted numerous research efforts focusing on road safety. PANDA You et al. (2022) predicted road risks after natural disasters by combining trajectories data and road event data. To identify road obstacles such as fallen trees and ponding water, RADAR Chen et al. (2018) utilized co-training and active learning to fuse the heterogeneous features from trajectory data and environmental sensing data. Yin et al. (2023b) proposed a multi-modal fusion network for inferencing missing road attributes. In their methodology (shown in Figure 22), the robustness of the network is largely enhanced by the pixel-level fusion of GPS traces and satellite images.
@@
Figure 22: Network architecture of the proposed multi-modal fusion network by Yin et al. (2023b) for robust road attribute detection.
@@
With the rapid urbanization of numerous countries, abnormal traffic incidents, have emerged as a substantial threat to public health and development. Wang et al. (2021e) used a Multi-View Multi-Task Spatio-Temporal Network to capture the various context features such as weather, POIs, and road conditions, which result in the occurrence of traffic accidents. The multi-task learning framework is effective in modeling features of different granularity levels, thanks to the tie of spatial associations.
@@
Human factors play a critical role in transportation safety and can be considered a primary factor in accidents Negash and Yang (2023); Mozaffari et al. (2020). SAX-DF Liu et al. (2022b) is a Symbolic Aggregate approximation Data Fusion model to comprehensively utilize multi-source and heterogeneous data, which can effectively improve the driver behavior detection performance with the help of a Positive Danger Mapping algorithm. The rise of autonomous driving has stimulated the demand for real-time driver behavior detection. Huang et al. (2022b) proposed a deformable inverted residual network to adaptively detect real-time driver behavior.
@@
Flow control is an important downstream application that holds great significance for public policies such as resource scheduling and urban planning Gerla and Kleinrock (1980). For instance, utilizing urban flow data enables governments to implement flow control measures during events such as New Year’s Eve at Shanghai’s Bund or Times Square in New York. By managing the crowd flow and directing subway passengers to nearby stations, the government can prevent potential dangers and promote efficient social functioning, thus yielding social benefits. Private ride-hailing platforms like DiDi and Uber allocate more transportation resources to densely populated areas, which allows them to gain better economic benefits.
@@
Indeed, flow control can be broadly categorized into two main areas: i) Traffic Flow Control Bellemans et al. (2002), which involves regulating and optimizing the flow of vehicles on roads and highways. ii) Mobility Prediction Zhang and Dai (2018); Barbosa et al. (2018), which endeavors to forecast and comprehend patterns of mobility behavior exhibited by humans or vehicles, is intended to enhance the planning and management of urban transportation and resources.
@@
DeepSTN+ Lin et al. (2019) utilized a spatio-temporal network to predict inflow and outflow regarding POIs and historical data. To infer urban flow, UrbanSTC Qu et al. (2022) implemented contrastive pretraining in both spatial and temporal to learn robust features in both two modalities with different pretasks. Likewise, CSST Ke et al. (2023) made use of a pretrain-finetuning paradigm to align low-quality GPS reports and external factors with the crowd flow.
@@
For mobility prediction, Miyazawa et al. (2019) combined GPS traces and geo-tagged tweets to model human crowd flow. GraphTUL Gao et al. (2022c) solved the Trajectory user linking (TUL) problem by training an adversarial network in a semi-supervised way.
@@
DMVST-Net Yao et al. (2018), shown in Figure 23, combined multi-view information to implement taxi demand prediction. It splits the process into spatial, temporal, and semantic views, which are modeled by local CNN, LSTM, and semantic graph embedding, respectively. To address the challenge of predicting travel demands in city regions for future time intervals demands, DeepTP Yuan et al. (2021a) proposed to encode and capture three key properties from traffic data: Region-Level Correlations, Temporal Periodicity, and Inter-Traffic Correlations. To jointly predict demands for various ride-hailing service modes like solo and shared rides, Ke et al. (2021) combined multi-graph convolutional networks with two multi-task learning structures, enhancing prediction accuracy for diverse ride-hailing services in urban areas. For delivery tasks, MetaSTP Ruan et al. (2022) introduced a meta-learning-based neural network model for predicting service time in last-mile delivery. Utilizing a Transformer-based layer and location prior knowledge, MetaSTP addresses complex delivery scenarios, showing significant improvements in prediction accuracy and practical deployment in JD Logistics.
Figure 23: The architecture of DMVST-Net Yao et al. (2018), which is designed to incorporate spatial, temporal, and semantic components for taxi demand prediction.
The prediction of passenger demand is also of great significance, which can assist transportation departments and related enterprises in enhancing the planning and management of passenger traffic. Bai et al. (2019b) proposed a deep learning framework combining graph convolutional recurrent neural networks and LSTM networks. This framework aims to accurately predict citywide passenger demand in ride-sharing platforms by analyzing historical demand data and external factors like weather and time. STG2Seq Bai et al. (2019a) solved the challenges for multi-step passenger demand forecasting in cities. It combines a Graph Convolutional Network (GCN) with an innovative encoder and attention-based output module, effectively capturing spatio-temporal correlations in-demand data and outperforming traditional methods in real-world tests.
@@ 5.2.2 Transportation Recommendation
Figure 24: An overview of unified route representation learning for multi-modal transportation recommendation Liu et al. (2022a).
@@
Different from public services, transportation recommendations are proposed for specific personalized needs, significantly facilitating human society’s daily life, thus receiving increasing attention in recent years. CondorFerries (2023) demonstrated that the prevalence of private travel has been on the rise since 2016, with 83 million Americans looking to take a solo trip in 2023. With the progress of human society and the development of urbanization, transportation recommendations are expected to create more economic and social benefits in the future. Among the various applications of transportation recommendations, route Recommendations, multi-modal transportation recommendations, and trip recommendations have garnered the most research attention. In this section, we will primarily introduce them respectively.
@@
Route recommendation plays a core role in many applications such as taxi services like RoD (ride-on-demand) and navigation Li et al. (2017a); Dai et al. (2015). Guo et al. (2022) presented a novel force-directed algorithm for improving route recommendations in ride-on-demand services. This method, inspired by electrostatic principles, utilizes urban data to align vehicle distribution with passenger demand, thereby enhancing route efficiency.
@@
Given the origin and destination, multi-modal transportation recommendations offer travel plans consisting of diverse transportation modes (e.g., driving, cycling, and public transit), as well as instructions for making seamless transitions between different modes. Doing so improves individuals’ convenience and well-being, particularly for older adults and children. Liu et al. (2019b) proposed Trans2Vec, a model that enhances transportation mode recommendations by integrating heterogeneous data to capture user and location preferences, thereby improving accuracy in suggesting multi-modal transportation options. Liu et al. (2022a) developed a novel framework (shown in Figure 24 for multi-modal transportation recommendations. The approach leverages a hierarchical multi-task route representation learning technique that integrates spatio-temporal autocorrelation modeling and coherent-aware attentive route learning. Enhanced by spatio-temporal pre-training strategies, the framework effectively utilizes time-dependent multi-view transportation graphs, offering a more nuanced and accurate transportation recommendation system.
@@
Trip Recommendation aims to propose a series of POIs for individuals with specific travel preferences, such as the number of attractions, starting and ending points, and so on. He et al. (2019) introduced a novel context-aware POI embedding model that jointly learns the impact of POI popularity, user preferences, and co-occurring POIs. Uniquely integrating visual features, Photo2Trip Zhao et al. (2017) utilized geo-tagged photos with collaborative filtering models to enhance personalized tour recommendations. GraphTrip Gao et al. (2023a) addressed the sparsity of data and the need for more personalized trip recommendations by effectively integrating diverse knowledge domains. The paper introduces a dual-grained mobility learning approach that uses spatio-temporal graph representation.
@@ 5.3 Economy
Real estate price is a vital indicator for a city which is affected by various factors. To manage citywide real estate information, Du et al. (2019) designed a dual-level collective learning (DLCL) framework to collectively learn spatial representation through intra-region and inter-region geographical structure knowledge which contain POI information as well as geographical and human mobility information for accurate prediction of the real estate price of Beijing. In their framework, intra-region POI and other side information are constructed through multi-view POI-POI networks. The constructed intra-region structure and inter-region autocorrelation are embedded through the Adversarial AutoEncoder.
Jenkins et al. (2019) proposed a model named RegionEncoder for various data sources and conducted experiments on Chicago and New York City for the perdition of real estate prices as well. They added satellite image data to the input of the model as a supplementary and gained impressive results in the experiment.
In addition to real estate price prediction, the study conducted by Xi et al. (2022) introduced an attention model that combines satellite imagery with POI information to estimate various economic indicators. Their methodology addresses the limitations of previous studies on satellite image data by leveraging the complementary nature of POI data in capturing human activity factors. Li et al. (2022b) conducted the first research to organize multi-view urban images (including satellite images and street-view images) by leveraging city structural information. The effective combination of the street-view image data contributed to a significant improvement of 10% in urban economic indicators predicting compared to previous studies before 2022. Besides, Liu et al. (2023g) also emphasized the necessity of integrating street-view images with satellite images in their research about urban economy.
@@ 5.4 Public Safety and Security
In the process of daily urban operations, ensuring public safety and security is a critical responsibility of cities and a standing focus for city administrators and researchers. The activities of humans and the movement of vehicles often encounter various safety hazards. Additionally, natural events like landslides and air pollution can also pose threats to people’s lives. The various urban big data provide us with opportunities to gain insights into unsafe factors and predict unsafe events. By integrating and analyzing diverse data sources such as human mobility, social media data, sensor data, and emergency response records, we can identify patterns and trends that may indicate potential safety risks. For example, Zhang et al. (2021b) proposed a multi-view region embedding framework with human mobility, region attribution, and crime records of New York City to predict the number of crime events in each region. Jiang et al. (2021) designed a framework to utilize vehicle trajectories and road environment data to infer traffic violation-prone locations in Xiamen City.
Traffic safety, being a paramount concern for numerous countries, has emerged as a prominent focus in recent data fusion studies pertaining to urban security. Wang et al. (2021a) proposed a Geographical and Semantic spatio-temporal Network (GSNet) model for predicting traffic accidents from taxi order, POI, and weather data. They conducted an experiment on New York City and Chicago based on real-world traffic accident datasets, the result demonstrated a satisfactory result for their model. Then Wang et al. (2023c) designed a cross-scale feature fusion mechanism to integrate features from different scale data as well as a feature fusion component to integrate multi-view features. Based on this work, a multi-task learning framework that can simultaneously predict fine and coarse-grained traffic accident risks was proposed.
Environment safety and the mitigation of natural disasters have also been remarkable research focal points. These areas of study have garnered significant attention due to their profound implications for urban environments and the well-being of communities. Researchers have made notable strides in leveraging data fusion techniques to enhance environmental monitoring, early warning systems, and disaster response strategies. For instance, based on the existing study of trajectories, Luo et al. (2021a) combined traffic trajectory information with road networks to efficiently identify traffic bottlenecks on the road. Chen et al. (2018) and You et al. (2022) focused on studying the road obstacle situation or road risks after disasters in a city from its vehicle trajectory, satellite image, and meteorology data. Besides, Song et al. (2017) built an intelligent system named DeepMob in Japan based on users trajectories, earthquake records, text reports, and transportation network data to analyze and predict human behavior and mobility following natural disasters.
@@ 5.5 Social
The advances in wireless communication and location acquisition technologies enable people to add a location dimension to traditional social networks Narayanan and Cherukuri (2016). With the advancement of data fusion advantage and the popularity of recommendation devices, services based on Location-based Social Networks (LBSN) have penetrated people’s lives. Users on social media generate a variety of content with geo-information every day. Such posts combine geographical coordinates (always in GPS location) and user-generated content (text, image, or audio) which might be associated with the semantic meaning of those places. It is strongly promising to fusion the geographical information and other information through the users’ posts to bridge the gap between users’ activities in digital and physical worlds and contribute to various downstream tasks such us POI or friend recommendation Chen et al. (2015); Cao et al. (2020); Park and Han (2018); Fang et al. (2023); Li et al. (2023c) and community analysis and detection Talpur and Zhang (2018); Liu et al. (2022c); Xu et al. (2023a).
Recommendation on LBSN based on deep learning is a complex task in real applications as the irregular spatial structure of geo-location data is non-Euclidean and traditional neural network-based deep learning is incapable of such non-Euclidean data. So graph neural network is widely used to understand the spatial structure information with geo-location data. Fang et al. (2023) proposed a multi-graph fusion approach for POI recommendation by constructing a user-POI interaction graph on LBSN. For LBSN friend recommendation, Li et al. (2023c) designed a learning framework by leveraging multi-graph to model raw LBSN data and defining various connections between nodes to represent spatio-temporal information. In their methodology, a contrastive learning model was proposed to integrate spatio-temporal features of human trajectories in cities for user node embedding learning.

### v2

@@ 5.2 Transportation
Transportation serves as the arteries of a city, acting as bridges that connect different urban entities [62]. As a result, the comprehensive utilization of multi-modal data focuses a significant proportion of attention on downstream transportation-related tasks [199, 27, 245, 148].
@@ 5.2.1 Public Transportation Service
The advancement of a city is primarily demonstrated by enhancements in its transportation systems [62, 91, 52]. Public transportation, in comparison to private transportation, is recognized for its superior environmental sustainability and energy efficiency [199]. Guided by government supervision, public transportation embodies the features of public goods. In addition to fundamental passenger transportation, there are also freight transportation services, along with emerging services such as food delivery and ride-hailing, all of which contribute to enhancing urban convenience. Different from [391], where transportation systems were categorized based on the type of vehicles used, this classification of services is based on their specific application scenarios. Public transportation applications can primarily be categorized into three main groups: Safety, Flow Control, and Efficiency.
@@
Safety is the most fundamental and urgent requirement. Consequently, roads, being high-risk areas for public transportation safety, have prompted numerous research efforts focusing on road safety. PANDA [342] predicted road risks after natural disasters by combining trajectories data and road event data. To identify road obstacles such as fallen trees and ponding water, RADAR [39] utilized co-training and active learning to fuse the heterogeneous features from trajectory data and environmental sensing data. Yin et al. [337] proposed a multi-modal fusion network for inferencing missing road attributes. In their methodology (shown in Figure 24), the robustness of the network is largely enhanced by the pixel-level fusion of GPS traces and satellite images.
@@
Figure 24: Network architecture of the proposed multi-modal fusion network by Yin et al. [337] for robust road attribute detection.
@@
With the rapid urbanization of numerous countries, abnormal traffic incidents, have emerged as a substantial threat to public health and development. [287] used a Multi-View Multi-Task Spatio-Temporal Network to capture the various context features such as weather, POIs, and road conditions, which result in the occurrence of traffic accidents. The multi-task learning framework is effective in modeling features of different granularity levels, thanks to the tie of spatial associations.
@@
Human factors play a critical role in transportation safety and can be considered a primary factor in accidents [205, 201]. SAX-DF [178] is a Symbolic Aggregate approximation Data Fusion model to comprehensively utilize multi-source and heterogeneous data, which can effectively improve the driver behavior detection performance with the help of a Positive Danger Mapping algorithm. The rise of autonomous driving has stimulated the demand for real-time driver behavior detection. Huang et al. [111] proposed a deformable inverted residual network to adaptively detect real-time driver behavior.
@@
Flow control is an important downstream application that holds great significance for public policies such as resource scheduling and urban planning [86]. For instance, utilizing urban flow data enables governments to implement flow control measures during events such as New Year’s Eve at Shanghai’s Bund or Times Square in New York. By managing the crowd flow and directing subway passengers to nearby stations, the government can prevent potential dangers and promote efficient social functioning, thus yielding social benefits. Private ride-hailing platforms like DiDi and Uber allocate more transportation resources to densely populated areas, which allows them to gain better economic benefits.
@@
Indeed, flow control can be broadly categorized into two main areas: i) Traffic Flow Control [17], which involves regulating and optimizing the flow of vehicles on roads and highways. ii) Mobility Prediction [360, 15], which endeavors to forecast and comprehend patterns of mobility behavior exhibited by humans or vehicles, is intended to enhance the planning and management of urban transportation and resources.
@@
DeepSTN+ [170] utilized a spatio-temporal network to predict inflow and outflow regarding POIs and historical data. To infer urban flow, UrbanSTC [226] implemented contrastive pretraining in both spatial and temporal to learn robust features in both two modalities with different pretasks. Likewise, CSST [133] made use of a pretrain-finetuning paradigm to align low-quality GPS reports and external factors with the crowd flow.
@@
For mobility prediction, [200] combined GPS traces and geo-tagged tweets to model human crowd flow. GraphTUL [79] solved the Trajectory user linking (TUL) problem by training an adversarial network in a semi-supervised way.
@@
DMVST-Net [329], shown in Figure 25, combined multi-view information to implement taxi demand prediction. It splits the process into spatial, temporal, and semantic views, which are modeled by local CNN, LSTM, and semantic graph embedding, respectively. To address the challenge of predicting travel demands in city regions for future time intervals demands, DeepTP [349] proposed to encode and capture three key properties from traffic data: Region-Level Correlations, Temporal Periodicity, and Inter-Traffic Correlations. To jointly predict demands for various ride-hailing service modes like solo and shared rides, Ke et al. [132] combined multi-graph convolutional networks with two multi-task learning structures, enhancing prediction accuracy for diverse ride-hailing services in urban areas. For delivery tasks, MetaSTP [235] introduced a meta-learning-based neural network model for predicting service time in last-mile delivery. Utilizing a Transformer-based layer and location prior knowledge, MetaSTP addresses complex delivery scenarios, showing significant improvements in prediction accuracy and practical deployment in JD Logistics.
Figure 25: The architecture of DMVST-Net [329], which is designed to incorporate spatial, temporal, and semantic components for taxi demand prediction.
The prediction of passenger demand is also of great significance, which can assist transportation departments and related enterprises in enhancing the planning and management of passenger traffic. Bai et al. [12] proposed a deep learning framework combining graph convolutional recurrent neural networks and LSTM networks. This framework aims to accurately predict citywide passenger demand in ride-sharing platforms by analyzing historical demand data and external factors like weather and time. STG2Seq [11] solved the challenges for multi-step passenger demand forecasting in cities. It combines a Graph Convolutional Network (GCN) with an innovative encoder and attention-based output module, effectively capturing spatio-temporal correlations in-demand data and outperforming traditional methods in real-world tests.
@@ 5.2.2 Transportation Recommendation
Figure 26: An overview of unified route representation learning for multi-modal transportation recommendation [173].
@@
Different from public services, transportation recommendations are proposed for specific personalized needs, significantly facilitating human society’s daily life, thus receiving increasing attention in recent years. CondorFerries [48] demonstrated that the prevalence of private travel has been on the rise since 2016, with 83 million Americans looking to take a solo trip in 2023. With the progress of human society and the development of urbanization, transportation recommendations are expected to create more economic and social benefits in the future. Among the various applications of transportation recommendations, route Recommendations, multi-modal transportation recommendations, and trip recommendations have garnered the most research attention. In this section, we will primarily introduce them respectively.
@@
Route recommendation plays a core role in many applications such as taxi services like RoD (ride-on-demand) and navigation [157, 52]. Guo et al. [93] presented a novel force-directed algorithm for improving route recommendations in ride-on-demand services. This method, inspired by electrostatic principles, utilizes urban data to align vehicle distribution with passenger demand, thereby enhancing route efficiency.
@@
Given the origin and destination, multi-modal transportation recommendations offer travel plans consisting of diverse transportation modes (e.g., driving, cycling, and public transit), as well as instructions for making seamless transitions between different modes. Doing so improves individuals’ convenience and well-being, particularly for older adults and children. Liu et al. [176] proposed Trans2Vec, a model that enhances transportation mode recommendations by integrating heterogeneous data to capture user and location preferences, thereby improving accuracy in suggesting multi-modal transportation options. Liu et al. [173] developed a novel framework (shown in Figure 26 for multi-modal transportation recommendations. The approach leverages a hierarchical multi-task route representation learning technique that integrates spatio-temporal autocorrelation modeling and coherent-aware attentive route learning. Enhanced by spatio-temporal pre-training strategies, the framework effectively utilizes time-dependent multi-view transportation graphs, offering a more nuanced and accurate transportation recommendation system.
@@
Trip Recommendation aims to propose a series of POIs for individuals with specific travel preferences, such as the number of attractions, starting and ending points, and so on. He et al. [99] introduced a novel context-aware POI embedding model that jointly learns the impact of POI popularity, user preferences, and co-occurring POIs. Uniquely integrating visual features, Photo2Trip [386] utilized geo-tagged photos with collaborative filtering models to enhance personalized tour recommendations. GraphTrip [77] addressed the sparsity of data and the need for more personalized trip recommendations by effectively integrating diverse knowledge domains. The paper introduces a dual-grained mobility learning approach that uses spatio-temporal graph representation.
@@ 5.3 Economy
Real estate price is a vital indicator for a city which is affected by various factors. To manage citywide real estate information, Du et al. [63] designed a dual-level collective learning (DLCL) framework to collectively learn spatial representation through intra-region and inter-region geographical structure knowledge which contain POI information as well as geographical and human mobility information for accurate prediction of the real estate price of Beijing. In their framework, intra-region POI and other side information are constructed through multi-view POI-POI networks. The constructed intra-region structure and inter-region autocorrelation are embedded through the Adversarial AutoEncoder.
Jenkins et al. [116] proposed a model named RegionEncoder for various data sources and conducted experiments on Chicago and New York City for the perdition of real estate prices as well. They added satellite image data to the input of the model as a supplementary and gained impressive results in the experiment.
In addition to real estate price prediction, the study conducted by Xi et al. [309] introduced an attention model that combines satellite imagery with POI information to estimate various economic indicators. Their methodology addresses the limitations of previous studies on satellite image data by leveraging the complementary nature of POI data in capturing human activity factors. Li et al. [154] conducted the first research to organize multi-view urban images (including satellite images and street-view images) by leveraging city structural information. The effective combination of the street-view image data contributed to a significant improvement of 10% in urban economic indicators predicting compared to previous studies before 2022. Besides, Liu et al. [186] also emphasized the necessity of integrating street-view images with satellite images in their research about urban economy.
@@ 5.4 Public Safety and Security
In the process of daily urban operations, ensuring public safety and security is a critical responsibility of cities and a standing focus for city administrators and researchers. The activities of humans and the movement of vehicles often encounter various safety hazards. Additionally, natural events like landslides and air pollution can also pose threats to people’s lives. The various urban big data provide us with opportunities to gain insights into unsafe factors and predict unsafe events. By integrating and analyzing diverse data sources such as human mobility, social media data, sensor data, and emergency response records, we can identify patterns and trends that may indicate potential safety risks. For example, Zhang et al. [368] proposed a multi-view region embedding framework with human mobility, region attribution, and crime records of New York City to predict the number of crime events in each region. Jiang et al. [124] designed a framework to utilize vehicle trajectories and road environment data to infer traffic violation-prone locations in Xiamen City.
Traffic safety, being a paramount concern for numerous countries, has emerged as a prominent focus in recent data fusion studies pertaining to urban security. Wang et al. [272] proposed a Geographical and Semantic spatio-temporal Network (GSNet) model for predicting traffic accidents from taxi order, POI, and weather data. They conducted an experiment on New York City and Chicago based on real-world traffic accident datasets, the result demonstrated a satisfactory result for their model. Then Wang et al. [288] designed a cross-scale feature fusion mechanism to integrate features from different scale data as well as a feature fusion component to integrate multi-view features. Based on this work, a multi-task learning framework that can simultaneously predict fine and coarse-grained traffic accident risks was proposed.
Environment safety and the mitigation of natural disasters have also been remarkable research focal points. These areas of study have garnered significant attention due to their profound implications for urban environments and the well-being of communities. Researchers have made notable strides in leveraging data fusion techniques to enhance environmental monitoring, early warning systems, and disaster response strategies. For instance, based on the existing study of trajectories, Luo et al. [188] combined traffic trajectory information with road networks to efficiently identify traffic bottlenecks on the road. Chen et al. [39] and You et al. [342] focused on studying the road obstacle situation or road risks after disasters in a city from its vehicle trajectory, satellite image, and meteorology data. Besides, Song et al. [249] built an intelligent system named DeepMob in Japan based on users trajectories, earthquake records, text reports, and transportation network data to analyze and predict human behavior and mobility following natural disasters.
@@ 5.5 Social
The advances in wireless communication and location acquisition technologies enable people to add a location dimension to traditional social networks [204]. With the advancement of data fusion advantage and the popularity of recommendation devices, services based on Location-based Social Networks (LBSN) have penetrated people’s lives. Users on social media generate a variety of content with geo-information every day. Such posts combine geographical coordinates (always in GPS location) and user-generated content (text, image, or audio) which might be associated with the semantic meaning of those places. It is strongly promising to fusion the geographical information and other information through the users’ posts to bridge the gap between users’ activities in digital and physical worlds and contribute to various downstream tasks such us POI or friend recommendation [42, 31, 219, 69, 155] and community analysis and detection [256, 184, 316].
Recommendation on LBSN based on deep learning is a complex task in real applications as the irregular spatial structure of geo-location data is non-Euclidean and traditional neural network-based deep learning is incapable of such non-Euclidean data. So graph neural network is widely used to understand the spatial structure information with geo-location data. Fang et al. [69] proposed a multi-graph fusion approach for POI recommendation by constructing a user-POI interaction graph on LBSN. For LBSN friend recommendation, Li et al. [155] designed a learning framework by leveraging multi-graph to model raw LBSN data and defining various connections between nodes to represent spatio-temporal information. In their methodology, a contrastive learning model was proposed to integrate spatio-temporal features of human trajectories in cities for user node embedding learning.